OpenStudy (anonymous):

16100 people at a concert 2/5 of the audience are boys, 3/7 of the audience are girls, How many girls are there? What fraction of the people are adults?

OpenStudy (luffingsails):

Hi Jake, So, the total number of people is given by: Number of Boys + Number of Girls + Number of Adults = 16100. We know: (2/5)*16100 = Number of Boys (3/7)*16100 = Number of Girls So, we can place those into the equation. (2/5)*16100 + (3/7)*16100 + Number of Adults = 16100 Now just solve for Number of adults and you are done. Help enough?

OpenStudy (anonymous):

Yes thats great but how many girls are there?

OpenStudy (luffingsails):

The Number of girls is given by (3/7)*16100.

OpenStudy (luffingsails):

Sorry. So, for the fraction of adults... Just do: 1 - (2/5)-(3/7) = Fraction of adults.

OpenStudy (anonymous):

rightthanks!:)

OpenStudy (luffingsails):

Cool. Sorry for the earlier stuff. I'll read the question better next time. <laugh>

OpenStudy (anonymous):

no problem:)